OpenStudy (anonymous):

This is false right?: 4x2 + 4x + 1 = (2x + 1)2

OpenStudy (dape):

Nope, that holds for all x.

OpenStudy (dape):

Proof: \[ (2x+1)^2=(2x+1)*(2x+1)=2x*2x+2x*1+1*2x+1*1 \\ = 4x^2+4x+1 \]

OpenStudy (anonymous):

Thnk u.